OpenStudy (anonymous):

Solve the system using any method. 1)3x+5y=-1 2x-y=-5

OpenStudy (amistre64):

any method?

OpenStudy (anonymous):

substion

OpenStudy (amistre64):

x y 3 5 = -1 2 -1 =-5 y = (-15+2)/(-3-10) x = (-25+1)/(10+3)

OpenStudy (anonymous):

spell wrong sorry

OpenStudy (anonymous):

so whats y and whats x

OpenStudy (amistre64):

well, when you have just numbers, you tend to work thru the math like they did back in kindergarten :) y = (-15+2)/(-3-10) = -13/-13 =1 x = (-25-1)/(10+3) = -26/13 =-2

OpenStudy (anonymous):

y =1 and x=-2

OpenStudy (amistre64):

yep; that method is known as cramers method, i figured it out before i know that there was a method for it ....

OpenStudy (anonymous):

okay thanks so thats works as the subsatotion method also

OpenStudy (amistre64):

yes; it tends to be the quickest method since it just follows a pattern that pulls the numbers out rather then alot of mental mathematical gymnastics

OpenStudy (anonymous):

okay thanks do you think you couldl help me with a different problem

OpenStudy (anonymous):

Tell whether this given ordered triple is a solution of a system. (3,3,4) X+2y-2z=1 7x-4y+z=11 2x-3y+2z=5

OpenStudy (amistre64):

given 2 equations: ax + by = n and cx + dy = m \begin{array}c a&b&n\\ c&d&m\\\\\end{array}\[y=\frac{ma-nc}{ad-bc}\]\[x=\frac{mb-nd}{bc-ad}\]

OpenStudy (amistre64):

the only way to tell is to plug them in to see

OpenStudy (amistre64):

25 -12 ----- 13 not 11

OpenStudy (anonymous):

i am lost

OpenStudy (anonymous):

so no its not a solution

OpenStudy (amistre64):

it dont fit the 2nd equation, so it cant be a solution